# ISRE-0001-2024.00

### v1

@@ 4.2.2 Results and Analysis
Following the procedure detailed in Algorithm 1, we apply DeCoLe to the dataset introduced in Section 4.2.1 to perform bias-aware mislabeling detection. We compare the performance of DeCoLe with those of two state-of-the-art alternative frameworks: Confident Learning (CL) (Northcutt et al. 2021) and Co-Teaching (CoT) (Han et al. 2018), along with random sampling which provides a useful reference. CL is the most directly related method, as it operates on the same core principles as DeCoLe but assumes noise to be solely class dependent. Co-Teaching, designed for robust learning under high noise, trains two networks in parallel, each selecting small-loss (likely clean) examples to train its peer. To adapt it for mislabeling detection, we use the discarded large-loss instances as its predicted mislabeled set. We use logistic regression as the base model (classifier) for all approaches. We compute 95% confidence intervals by performing 20 iterations with different random seeds for data shuffling and model initialization to ensure the robustness and consistency of performance evaluation.
@@
We present the results for model performance across different algorithms and demographic groups in Figure 4, using the performance measures defined in Section 4.1. Specifically, Figure  4 (a) examines the overall mislabeling detection performance, where the x-axis, $\textit{Recall}_{\bm{\hat{D}}_{\tilde{y}\neq y^{*}|g}}$, quantifies each model’s ability to recall mislabeled instances, while the y-axis, $\textit{Precision}_{\bm{\hat{D}}_{\tilde{y}=y^{*}|g}}$, reflects the precision of observed labels in data estimated as correctly labeled. Models positioned further to the right demonstrate stronger recall, while higher placement on the y-axis indicates greater precision. DeCoLe consistently occupies the upper-right region of the plots, achieving both the highest recall of mislabeled instances and the highest label precision in data estimated as correclty labeled.
Figure 4 (b) further examines the bias-aware aspect of mislabeling detection by plotting the recall of bias-inducing errors, $\textit{Recall}_{\bm{\hat{D}}_{\tilde{y}=c,y^{*}=\overline{c}|g}}$, on the x-axis and the precision of the bias-dominant class in the data estimated as correclty labeled, $\textit{Precision}_{\bm{\hat{D}}_{\tilde{y}=y^{*}=c|g}}$ on the y-axis. A position further to the right reflects higher recall of bias-inducing errors, while a higher position indicates greater precision for the bias-dominant class. DeCoLe consistently appears in the upper-right area of the plots, indicating both the highest recall of bias-inducing errors and the highest precision for the bias-dominant class in the data estimated as correctly labeled.
@@
Table 2: Overall recall of mislabeled instances ($\textit{Recall}_{\bm{\hat{D}}_{\tilde{y}\neq y^{*}|g}}$) using DeCoLe versus competing algorithms. Results are shown for simulations under different error rates, and are disaggregated by demographic groups. Performance values are reported with 95% confidence intervals (± value). Bold denotes the best-performing model, (*) denotes statistically significant improvements over the second-best algorithm. DeCoLe consistently outperforms competing algorithms.
@@
Table 3: Overall precision of the observed label in data estimated as correctly labeled ($\textit{Precision}_{\bm{\hat{D}}_{\tilde{y}=y^{*}|g}}$) using DeCoLe versus competing algorithms under different error rates. Results are disaggregated by demographic groups. Performance values are reported with 95% confidence intervals (± value). Bold denotes the best-performing model, while an asterisk (*) denotes statistically significant improvements over the second-best algorithm. DeCoLe consistently outperforms competing algorithms.
@@
Table 4: Recall of bias-inducing errors ($\textit{Recall}_{\bm{\hat{D}}_{\tilde{y}=c,y^{*}=\overline{c}|g}}$) using DeCoLe versus competing algorithms under different error rates. Results are disaggregated by demographic groups. Performance values are reported with 95% confidence intervals (± value). Bold denotes the best-performing model, while an asterisk (*) denotes statistically significant improvements over the second-best algorithm. DeCoLe consistently outperforms competing algorithms.
@@
Table 5: Precision of the bias-dominant class $\textit{Precision}_{\bm{\hat{D}}_{\tilde{y}=y^{*}=c|g}}$ across competing algorithms under different error rates. Results are disaggregated by different demographic groups. Performance values are reported with 95% confidence intervals (± value). Bold denotes the best-performing model, while an asterisk (*) denotes statistically significant improvements over competing methods. DeCoLe consistently outperforms competing algorithms.
@@
We further test DeCoLe in two more extended scenarios: First, we relax the assumption that $g_{0}$ is the minority group by simulating a balanced group setting where $g_{0}$ and $g_{1}$ each make up 50% of the population. We maintain asymmetric error rates such that $\pi_{0,{g_{0}}}>\pi_{1,{g_{1}}}$, ensuring that $g_{0}$ remains the group most affected by label bias, and we vary the severity of the errors. Second, we relax the assumption of asymmetric noise ($\pi_{0,{g_{0}}}>\pi_{1,{g_{1}}}$) by simulating a condition in which both groups experience the same error rate ($\pi_{0,{g_{0}}}=\pi_{1,{g_{1}}}$), tested at three error rate levels: 20%, 30%, and 40%. For this scenario, we consider $g_{0}$ to be a minority group that constitutes 30% of the population. In all cases, DeCoLe maintains superior performance across overall mislabeling detection (Metrics 1 & 2) and bias-aware performance (Metrics 3 & 4), reaffirming its robustness and fairness benefits. Detailed results for the extended settings can be found in Appendix 8.1 and 8.2.
@@ 4.3 Hate Speech Mislabeling Detection
Detecting and mitigating hate speech is an operational challenge for technology platforms (Jahan and Oussalah 2023). Meta, for example, defines hate speech as direct attacks on individuals based on protected characteristics—such as disability and religion—and enforces policies to remove dehumanizing speech or slurs that target individuals with these characteristics  (Meta Platforms, Inc. 2025). Given the vast volume of user-generated content produced online, and the detrimental psychological effects for human moderators who are exposed to disturbing content (Steiger et al. 2021), automated toxicity detection has been recognized as a scalable and cost-effective means to flag possible instances of hate speech  (Jahan and Oussalah 2023). However, the effectiveness of such systems hinges heavily on the quality of the data they rely on throughout the pipeline. When labels indicating whether a post is toxic are biased, analyses and models developed and evaluated on such data are at a higher risk of misclassifying content—either by failing to detect harmful speech or by incorrectly flagging benign posts (Sap et al. 2019, Davani et al. 2023). Leading platforms such as Meta and YouTube have invested significantly in labeling efforts and automated moderation pipelines. However, a persistent bottleneck remains: the potential bias embedded in the labels themselves (Vidgen et al. 2019). Likewise, labeled hate speech data has become a foundational resource in social science, enabling large-scale quantitative analyses of online behavior (Castaño-Pulgarín et al. 2021, Meza et al. 2018). As Castaño-Pulgarín et al. (2021) demonstrates, such data are routinely used to support both theoretical modeling and empirical validation in research on online hostility and identity-based violence. However, the presence of biased labels can compromise the integrity of these findings and lead to misguided policy insights. We apply DeCoLe to assess its effectiveness in large-scale bias-aware mislabeling detection as a means of addressing data integrity challenges in hate speech labeling.
@@
To rigorously evaluate mislabeling detection methods like DeCoLe, it is critical to have access to datasets that contain both noisy, real-world labels and higher-quality alternatives—a condition rarely met in practice. Fortunately, such an opportunity arises from the dataset introduced by Kennedy et al. (2020), which provides both commonly used hate speech labels ($\tilde{y}$) and theoretically grounded, more expensive labels ($y^{*}$), which are estimated on the basis of multiple dimensions of hate speech that are annotated separately, designed specifically to mitigate risks of error and bias in hate speech annotation. Specifically, to obtain a gold standard label $y^{*}$, Kennedy et al. (2020) proposes a novel approach based on Rasch Measurement Theory (RMT) to construct a higher quality hate speech label. Their measure articulates hate speech theoretically across eight dimensions (incite violence, humiliate, etc.), capturing the complexity of hate speech and limiting bias stemming from oversimplification and from idiosyncratic definitions of what constitutes hate speech. Furthermore, by evaluating inter-rater reliability, they are able to remove inconsistent raters, correcting human judgment biases and promoting reliability. The researchers assessed the validity and reliability of their proposed measurement approach and found that it demonstrates high internal consistency, test-retest reliability, and construct validity. In summary, Kennedy et al. (2020) limits bias in labels for hate speech, but it involves a significantly more costly labeling process, as it requires labels for each instance across eight different dimensions, rather than one. This precludes its widespread use in cost-sensitive settings. The dataset also has the advantage of collecting the more common labels used for hate speech, $\tilde{y}$, by directly asking if a post constitutes hate speech. Thus, the data contains a label $\tilde{y}$ and an improved label $y^{*}$ that we leverage to assess the performance of DeCoLe in a real-world setting from an impactful domain.
@@ 4.3.1 Experimental setting
To empirically evaluate DeCoLe’s ability in bias-aware mislabeling detection in hate speech data, we leverage the rare dual-label structure of the dataset introduced by Kennedy et al. (2020). Specifically, we treat the commonly used hate speech labels derived from a single survey item as observed labels ($\tilde{y}$), and use the theory-grounded, expert-curated annotations as the gold standard labels ($y^{*}$).
@@ 4.3.2 Empirical Results
Figure 5(a) displays overall performance, with the x-axis presenting the overall recall of mislabeled instances ($\textit{Recall}_{\bm{\hat{D}}_{\tilde{y}\neq y^{*}|g}}$) and the y-axis showing the overall precision of observed labels in data estimated to be correctly labeled ($\textit{Precision}_{\bm{\hat{D}}_{\tilde{y}=y^{*}|g}}$ ). Figure 5(b) presents bias-aware performance, plotting the recall of bias-inducing errors ($\textit{Recall}_{\bm{\hat{D}}{\tilde{y}=0,y^{*}=1|g}}$) on the x-axis and the precision of observed labels for the bias-dominant class in data estimated as correctly labeled ($\textit{Precision}_{\bm{\hat{D}}{\tilde{y}=y^{*}=0|g}}$) on the y-axis. In both subfigures, results are disaggregated by sexuality group, with each panel representing a distinct group. DeCoLe consistently appears in the upper-right portion of the plots across panels, indicating its superior effectiveness in mislabeling detection and its ability to do so in a bias-aware manner.
@@ 5.1.1 Building on DeCoLe in Data Quality Improvement
For data cleaning, as illustrated in our experiments, in the presence of labeling bias, pruning instances identified as mislabeled by DeCoLe can effectively reduce the rate of observed mislabeled instances relative to what can be achieved with the state-of-the-art alternatives. In the context of data relabeling, DeCoLe offers a versatile foundation for mitigating label bias across diverse contexts. Its fine-grained identification of mislabeled instances supports integration into human-in-the-loop workflows, enabling experts or labeling panels to focus review efforts on high-risk instances. This is particularly valuable in domains where acquiring high-quality labels is expensive or constrained by other resources such as time, allowing practitioners to allocate annotation resources more strategically. DeCoLe also lends itself to serve as a foundation for algorithmic-driven quality-enhancing approaches. For example, future work can explore its use within active relabeling frameworks that prioritize instances based on their likelihood of being mislabeled and the potential impact of correction on overall fairness (Bernhardt et al. 2022). These capabilities make DeCoLe well-suited not only for data cleaning and quality assurance but also for guiding strategically allocated expert relabeling efforts. By enabling such extensions, DeCoLe contributes not just a standalone solution but a flexible building block that opens a wide array of opportunities for advancing bias-aware data preparation.
@@ 5.2 Limitations and Future Works
Our empirical evaluation begins with synthetic data and is followed by real-world evidence in the context of hate speech detection. The use of one real-world dataset reflects a broader challenge in the field: the scarcity of datasets that simultaneously provide noisy observed labels, demographic attributes, and higher-quality labels necessary for the assessment of bias-aware mislabeling detection methods. While DeCoLe itself requires only observed labels and demographic attributes for application, effective evaluation demands access to a higher-quality reference—ideally a gold-standard label—to assess the extent of mislabeling and the effectiveness of its detection. The hate speech dataset introduced by Kennedy et al. (2020) offers a rare opportunity in this regard. Its theoretically grounded and more costly labeling process yields high-quality reference labels while also including the commonly used, noisier crowdsourced labels. Moreover, the dataset offers a rare opportunity to conduct evaluations across multiple demographic attributes—such as race, gender, and sexuality—each exhibiting unique linguistic and contextual patterns of hate speech. These differences give rise to distinct forms of label bias, enabling multiple evaluations of the method. Nonetheless, the broader lack of comparable datasets with dual-label structures and rich demographic annotations remains a critical limitation for advancing empirical research in bias-aware mislabeling detection. Future work would benefit from dedicated efforts to construct benchmark datasets for bias-aware mislabeling detection or any efforts facilitating label quality improvement.

### v2

@@ 4.2.2   Results and Analysis
Following the procedure detailed in Algorithm 1, we apply DeCoLe to the dataset introduced in Section 4.2.1 to perform bias-aware mislabeling detection. We compare the performance of DeCoLe with those of two state-of-the-art alternative frameworks: Confident Learning (CL) (northcutt2021confident, ) and Co-Teaching (CoT) (han2018co, ), along with random sampling which provides a useful reference. CL is the most directly related method, as it operates on the same core principles as DeCoLe but assumes noise to be solely class dependent. Co-Teaching, designed for robust learning under high noise, trains two networks in parallel, each selecting small-loss (likely clean) examples to train its peer. To adapt it for mislabeling detection, we use the discarded large-loss instances as its predicted mislabeled set. We use logistic regression as the base model (classifier) for all approaches. We compute 95% confidence intervals by performing 20 iterations with different random seeds for data shuffling and model initialization to ensure the robustness and consistency of performance evaluation.
@@
We present the results for model performance across different algorithms and demographic groups in Figure 4, using the performance measures defined in Section 4.1. Specifically, Figure  4 (a) examines the overall mislabeling detection performance, where the x-axis, $\textit{Recall}_{\boldsymbol{\hat{D}}_{\tilde{y}\neq y^{*}|g}}$, quantifies each model’s ability to recall mislabeled instances, while the y-axis, $\textit{Precision}_{\boldsymbol{\hat{D}}_{\tilde{y}=y^{*}|g}}$, reflects the precision of observed labels in data estimated as correctly labeled. Models positioned further to the right demonstrate stronger recall, while higher placement on the y-axis indicates greater precision. DeCoLe consistently occupies the upper-right region of the plots, achieving both the highest recall of mislabeled instances and the highest label precision in data estimated as correclty labeled.
Figure 4 (b) further examines the bias-aware aspect of mislabeling detection by plotting the recall of bias-inducing errors, $\textit{Recall}_{\boldsymbol{\hat{D}}_{\tilde{y}=c,y^{*}=\overline{c}|g}}$, on the x-axis and the precision of the bias-dominant class in the data estimated as correclty labeled, $\textit{Precision}_{\boldsymbol{\hat{D}}_{\tilde{y}=y^{*}=c|g}}$ on the y-axis. A position further to the right reflects higher recall of bias-inducing errors, while a higher position indicates greater precision for the bias-dominant class. DeCoLe consistently appears in the upper-right area of the plots, indicating both the highest recall of bias-inducing errors and the highest precision for the bias-dominant class in the data estimated as correctly labeled.
@@
Table 2: Overall recall of mislabeled instances ($\textit{Recall}_{\boldsymbol{\hat{D}}_{\tilde{y}\neq y^{*}|g}}$) using DeCoLe versus competing algorithms. Results are shown for simulations under different error rates, and are disaggregated by demographic groups. Performance values are reported with 95% confidence intervals (± value). Bold denotes the best-performing model, (*) denotes statistically significant improvements over the second-best algorithm. DeCoLe consistently outperforms competing algorithms.
@@
Table 3: Overall precision of the observed label in data estimated as correctly labeled ($\textit{Precision}_{\boldsymbol{\hat{D}}_{\tilde{y}=y^{*}|g}}$) using DeCoLe versus competing algorithms under different error rates. Results are disaggregated by demographic groups. Performance values are reported with 95% confidence intervals (± value). Bold denotes the best-performing model, while an asterisk (*) denotes statistically significant improvements over the second-best algorithm. DeCoLe consistently outperforms competing algorithms.
@@
Table 4: Recall of bias-inducing errors ($\textit{Recall}_{\boldsymbol{\hat{D}}_{\tilde{y}=c,y^{*}=\overline{c}|g}}$) using DeCoLe versus competing algorithms under different error rates. Results are disaggregated by demographic groups. Performance values are reported with 95% confidence intervals (± value). Bold denotes the best-performing model, while an asterisk (*) denotes statistically significant improvements over the second-best algorithm. DeCoLe consistently outperforms competing algorithms.
@@
Table 5: Precision of the bias-dominant class $\textit{Precision}_{\boldsymbol{\hat{D}}_{\tilde{y}=y^{*}=c|g}}$ across competing algorithms under different error rates. Results are disaggregated by different demographic groups. Performance values are reported with 95% confidence intervals (± value). Bold denotes the best-performing model, while an asterisk (*) denotes statistically significant improvements over competing methods. DeCoLe consistently outperforms competing algorithms.
@@
We further test DeCoLe in two more extended scenarios: First, we relax the assumption that $g_{0}$ is the minority group by simulating a balanced group setting where $g_{0}$ and $g_{1}$ each make up 50% of the population. We maintain asymmetric error rates such that $\pi_{0,{g_{0}}}>\pi_{1,{g_{1}}}$, ensuring that $g_{0}$ remains the group most affected by label bias, and we vary the severity of the errors. Second, we relax the assumption of asymmetric noise ($\pi_{0,{g_{0}}}>\pi_{1,{g_{1}}}$) by simulating a condition in which both groups experience the same error rate ($\pi_{0,{g_{0}}}=\pi_{1,{g_{1}}}$), tested at three error rate levels: 20%, 30%, and 40%. For this scenario, we consider $g_{0}$ to be a minority group that constitutes 30% of the population. In all cases, DeCoLe maintains superior performance across overall mislabeling detection (Metrics 1 & 2) and bias-aware performance (Metrics 3 & 4), reaffirming its robustness and fairness benefits. Detailed results for the extended settings can be found in Appendix C.1 and C.2.
@@ 4.3   Hate Speech Mislabeling Detection
Detecting and mitigating hate speech is an operational challenge for technology platforms (jahan2023systematic, ). Meta, for example, defines hate speech as direct attacks on individuals based on protected characteristics—such as disability and religion—and enforces policies to remove dehumanizing speech or slurs that target individuals with these characteristics  (metaHatefulConduct, ). Given the vast volume of user-generated content produced online, and the detrimental psychological effects for human moderators who are exposed to disturbing content (steiger2021, ), automated toxicity detection has been recognized as a scalable and cost-effective means to flag possible instances of hate speech  (jahan2023systematic, ). However, the effectiveness of such systems hinges heavily on the quality of the data they rely on throughout the pipeline. When labels indicating whether a post is toxic are biased, analyses and models developed and evaluated on such data are at a higher risk of misclassifying content—either by failing to detect harmful speech or by incorrectly flagging benign posts (sap2019risk, ; davani2023hate, ). Leading platforms such as Meta and YouTube have invested significantly in labeling efforts and automated moderation pipelines. However, a persistent bottleneck remains: the potential bias embedded in the labels themselves (vidgen2019challenges, ). Likewise, labeled hate speech data has become a foundational resource in social science, enabling large-scale quantitative analyses of online behavior (castano2021internet, ; meza2018targets, ). As castano2021internet demonstrates, such data are routinely used to support both theoretical modeling and empirical validation in research on online hostility and identity-based violence. However, the presence of biased labels can compromise the integrity of these findings and lead to misguided policy insights. We apply DeCoLe to assess its effectiveness in large-scale bias-aware mislabeling detection as a means of addressing data integrity challenges in hate speech labeling.
@@
To rigorously evaluate mislabeling detection methods like DeCoLe, it is critical to have access to datasets that contain both noisy, real-world labels and higher-quality alternatives—a condition rarely met in practice. Fortunately, such an opportunity arises from the dataset introduced by kennedy2020constructing , which provides both commonly used hate speech labels ($\tilde{y}$) and theoretically grounded, more expensive labels ($y^{*}$), which are estimated on the basis of multiple dimensions of hate speech that are annotated separately, designed specifically to mitigate risks of error and bias in hate speech annotation. Specifically, to obtain a gold standard label $y^{*}$, kennedy2020constructing proposes a novel approach based on Rasch Measurement Theory (RMT) to construct a higher quality hate speech label. Their measure articulates hate speech theoretically across eight dimensions (incite violence, humiliate, etc.), capturing the complexity of hate speech and limiting bias stemming from oversimplification and from idiosyncratic definitions of what constitutes hate speech. Furthermore, by evaluating inter-rater reliability, they are able to remove inconsistent raters, correcting human judgment biases and promoting reliability. The researchers assessed the validity and reliability of their proposed measurement approach and found that it demonstrates high internal consistency, test-retest reliability, and construct validity. In summary, kennedy2020constructing limits bias in labels for hate speech, but it involves a significantly more costly labeling process, as it requires labels for each instance across eight different dimensions, rather than one. This precludes its widespread use in cost-sensitive settings. The dataset also has the advantage of collecting the more common labels used for hate speech, $\tilde{y}$, by directly asking if a post constitutes hate speech. Thus, the data contains a label $\tilde{y}$ and an improved label $y^{*}$ that we leverage to assess the performance of DeCoLe in a real-world setting from an impactful domain.
@@ 4.3.1   Experimental setting
To empirically evaluate DeCoLe’s ability in bias-aware mislabeling detection in hate speech data, we leverage the rare dual-label structure of the dataset introduced by kennedy2020constructing . Specifically, we treat the commonly used hate speech labels derived from a single survey item as observed labels ($\tilde{y}$), and use the theory-grounded, expert-curated annotations as the gold standard labels ($y^{*}$).
@@ 4.3.2   Empirical Results
Figure 5(a) displays overall performance, with the x-axis presenting the overall recall of mislabeled instances ($\textit{Recall}_{\boldsymbol{\hat{D}}_{\tilde{y}\neq y^{*}|g}}$) and the y-axis showing the overall precision of observed labels in data estimated to be correctly labeled ($\textit{Precision}_{\boldsymbol{\hat{D}}_{\tilde{y}=y^{*}|g}}$ ). Figure 5(b) presents bias-aware performance, plotting the recall of bias-inducing errors ($\textit{Recall}_{\boldsymbol{\hat{D}}{\tilde{y}=0,y^{*}=1|g}}$) on the x-axis and the precision of observed labels for the bias-dominant class in data estimated as correctly labeled ($\textit{Precision}_{\boldsymbol{\hat{D}}{\tilde{y}=y^{*}=0|g}}$) on the y-axis. In both subfigures, results are disaggregated by sexuality group, with each panel representing a distinct group. DeCoLe consistently appears in the upper-right portion of the plots across panels, indicating its superior effectiveness in mislabeling detection and its ability to do so in a bias-aware manner.
@@ 5.1.1   Building on DeCoLe in Data Quality Improvement
For data cleaning, as illustrated in our experiments, in the presence of labeling bias, pruning instances identified as mislabeled by DeCoLe can effectively reduce the rate of observed mislabeled instances relative to what can be achieved with the state-of-the-art alternatives. In the context of data relabeling, DeCoLe offers a versatile foundation for mitigating label bias across diverse contexts. Its fine-grained identification of mislabeled instances supports integration into human-in-the-loop workflows, enabling experts or labeling panels to focus review efforts on high-risk instances. This is particularly valuable in domains where acquiring high-quality labels is expensive or constrained by other resources such as time, allowing practitioners to allocate annotation resources more strategically. DeCoLe also lends itself to serve as a foundation for algorithmic-driven quality-enhancing approaches. For example, future work can explore its use within active relabeling frameworks that prioritize instances based on their likelihood of being mislabeled and the potential impact of correction on overall fairness (bernhardt2022active, ). These capabilities make DeCoLe well-suited not only for data cleaning and quality assurance but also for guiding strategically allocated expert relabeling efforts. By enabling such extensions, DeCoLe contributes not just a standalone solution but a flexible building block that opens a wide array of opportunities for advancing bias-aware data preparation.
@@ 5.2   Limitations and Future Works
Our empirical evaluation begins with synthetic data and is followed by real-world evidence in the context of hate speech detection. The use of one real-world dataset reflects a broader challenge in the field: the scarcity of datasets that simultaneously provide noisy observed labels, demographic attributes, and higher-quality labels necessary for the assessment of bias-aware mislabeling detection methods. While DeCoLe itself requires only observed labels and demographic attributes for application, effective evaluation demands access to a higher-quality reference—ideally a gold-standard label—to assess the extent of mislabeling and the effectiveness of its detection. The hate speech dataset introduced by kennedy2020constructing offers a rare opportunity in this regard. Its theoretically grounded and more costly labeling process yields high-quality reference labels while also including the commonly used, noisier crowdsourced labels. Moreover, the dataset offers a rare opportunity to conduct evaluations across multiple demographic attributes—such as race, gender, and sexuality—each exhibiting unique linguistic and contextual patterns of hate speech. These differences give rise to distinct forms of label bias, enabling multiple evaluations of the method. Nonetheless, the broader lack of comparable datasets with dual-label structures and rich demographic annotations remains a critical limitation for advancing empirical research in bias-aware mislabeling detection. Future work would benefit from dedicated efforts to construct benchmark datasets for bias-aware mislabeling detection or any efforts facilitating label quality improvement.